Assume the economy can be described by the consumption function C = $300 billion + 0.9Y

1. Given the consumption function provided, the cumulative change in total spending if the government increases purchases of goods and services by $100 million will be:

a.

$100 million

b.

$80 million

c.

$ 5 billion

d.

$1 billion

2.

Given the consumption function provided, the cululative change in total spending if the government reduces taxes by $100 million will be:

a.

$90 million

b.

$900 million

c.

$50 million

d.

$100 million
Answer #1

1) MPC=change in C/change in Y=0.9

Govt. Spending multiplier=1/1-mpc=1/1-0.9=10

change in Aggregate spending=10*100=1000million=1billion.

ans is D

2)tax multiplier =-mpc/1-mpc=-0.9/1-0.9=-9

If taxes decreases by 100million then spending increases by 900million

ans is B